# An Empirical Study on LLM-based Agents for Automated Bug Fixing

### v1

@@ 4. Analysis & Results
We will sequentially present the analysis results and insights for RQ1 to RQ3.
@@ 4.1. RQ1: Effectiveness of Systems
We analyzed the versions of cases that each of the seven tools can solve individually, as well as the differences between the cases that each tool can resolve, as shown in Figure 1-(a). The histogram at the top of the figure shows the number of cases each tool can resolve in SWE-bench Lite. Specifically, ranked from lowest to highest, these are Agentless, Agentless+RepoGraph, AutoCodeRover, Alibaba Lingma Agent, Gru, Honeycomb, and MarsCode Agent, resolving 82, 89, 92, 99, 107, 115, and 118 cases, respectively. MarsCode Agent performs the best, achieving a 43.9% performance improvement over Agentless and addressing 39.3% of the total 300 cases in SWE-bench Lite. Compared to the popular APR benchmark Defects4J over the past decade, SWE-bench Lite introduces stricter usage protocols, prohibiting participants from leveraging dynamic evaluation results generated by closely related failing test cases as feedback information for filtering patches. This test case set can only be utilized as a quality standard once the patch generation process has concluded. In this context, many error localization methods based on dynamic test execution information—such as spectrum-based and mutation-based error localization methods—cannot be used, adding further complexity to problem detection and resolution. This strict protocol undoubtedly aligns more closely with real-world development scenarios, where repair tools must rely almost solely on issues raised by users and the current state of the code repository to devise solutions. Against this backdrop, MarsCode’s ability to address 39.3% of cases underscores its advanced capabilities and utility in real-world development environments.
(a) Issue Resolve Task
(b) File-level FL Task
(c) Line-level FL Task
Figure 1. Analysis of Performance of State-of-the-art Techniques on Issue Resolve and Fault Localization Tasks, where FL indicates Fault Localization.
Table 2. Issue Quality Metrics (REs indicates Reproducible Examples, and NL indicates Natutal Language).
Metrics Descriptions of Metrics Candidates Scores
Quality of Reproducible Examples This Metric indicates whether each issue description contains sufficient information to perform the desired task. Contains REs 10.00
Contains Partial REs 6.67
Info in NL 3.33
Not Enough Info 0.00
Quality of Resolve Solutions This metric indicates whether the solutions or steps to solve the problem are already provided in the issue description. Exact Patch 10.00
Complete Steps in NL 7.50
Some Steps in NL 5.00
No Solution 2.50
Misleading Information 0.00
Quality of Bug Locations (File/Function/Line) This metric indicates whether the issue description contains the correct bug location information. Stacktrace 10.00
Keyword 6.67
Natural Language 3.33
None Information 0.00
@@
In Figure 1-(a), the histogram beneath the tool names (referred to as Part-I) presents the case versions that different tool combinations can address. Each row indicates the number of case versions solvable by the tools marked with black dots but not by those marked with gray dots. For instance, in the first row, only MarsCode is marked with a black dot, while the other six tools are marked with gray dots, indicating that MarsCode can solve a unique set of 9 cases that none of the other six tools can handle. Similarly, the seventh row demonstrates that MarsCode and Honeycomb together can solve 5 cases that the other five tools cannot address. The final row shows that all seven tools collectively can solve 36 cases. The following sections will analyze the statistical findings presented in this figure from several perspectives.
Analysis of Case Solvability. Among the 300 cases in SWE-bench Lite, 168 cases are solvable by at least one of the seven tools (representing the sum of all values in Part-II of Figure 1-(a)), while 132 cases remain unsolved by any tool (not displayed in the figure and constituting the complement of the 168 solvable cases). Furthermore, 36 cases can be solved by all seven tools (represented by the last row in Part-II). We hypothesize that the issue descriptions for these 36 universally solvable cases are generally of higher quality, whereas the 132 cases that none of the tools can resolve likely exhibit lower-quality issue descriptions. To validate this hypothesis, we conducted a significance analysis of issue quality differences using five metrics provided by the issue quality analysis report in Agentless, as detailed in Table 2 (with bug location represented by three metrics: file, function, and line). For each metric, Agentless provided candidate values that represent the completeness of information, which we listed in descending order of completeness in the third column. We then assigned scores on a 10-point scale, uniformly distributing scores among candidate values to facilitate further analysis. For example, for the Quality of Reproducible Examples metric, we assigned a score of 10 points to "Contain REs" (highest information completeness), 6.67 points to "Contain Partial REs" (moderate completeness), 3.33 points to "Info in NL", and 0 points to "Not enough Info". While the number of candidates varied across metrics, we ensured that all metric scores ranged from 0 to 10, allowing comparability across metrics. Notably, for the Quality of Resolve Solutions metric, we positioned the "Misleading" candidate (where the issue description contains misleading information) at the lowest rank, as we consider it more detrimental than "No Solution." It is important to note that the indicators for each case have already been manually analyzed in the open-source artifacts of Agentless. Our approach builds on this by assigning scores to each candidate value to facilitate further analysis of issue quality.
@@
To examine the validity of our hypothesis, we calculated the mean values for the five indicators across two case sets: the set of 132 cases unsolved by any tool (referred to as the "No-one-resolve Set") and the set of 36 cases solvable by all tools (referred to as the "All-resolve Set"). These results are shown in Table 3. For each indicator, a higher score indicates that the corresponding issue description provides more complete and detailed information. It is evident that the scores for all five indicators are consistently higher in the All-resolve Set compared to the No-one-resolve Set. This suggests that issue quality significantly influences the effectiveness of resolution methods, underscoring the importance of crafting clear and comprehensive issue descriptions to improve resolution rates from the outset.
Examining the performance details for each indicator across both case sets, the most notable difference is observed in the Quality of Line-level Location metric, where the average score in the All-resolve Set is 27.8 times that of the No-one-resolve Set. This indicates that, at least within the SWE-bench Lite dataset, providing line-level location information markedly enhances the resolution rate for repair tools. The second largest difference appears in Quality of Solution in Description, with an average score difference of 2.13 times between the two sets, indicating that suggesting positive modifications based on observed symptoms can also substantially aid in improving tool solvability. This is followed by function-level and file-level location information, with score differences of 1.66 times and 1.47 times, respectively, suggesting that as the granularity of the bug location becomes coarser, its benefit to repair tools diminishes—a finding that aligns with intuitive expectations. Lastly, Quality of Reproducible Examples shows only a 1.1-time score difference between the two sets, suggesting that the completeness of this indicator has a comparatively minor impact on enhancing the solvability of repair tools.
In the above analysis, we validated our hypothesis by calculating differences across five metrics that represent issue quality in the All-resolve Set and No-one-resolve Set, showing a clear disparity in issue quality between cases that can be resolved and those that cannot. This reflects an overall trend, though some outliers may exist. On one hand, we observed several high-scoring cases in the No-one-resolve Set; we aim to investigate why, despite relatively complete information in these issues, no tools were able to resolve them. On the other hand, we also identified some low-scoring cases within the All-resolve Set, which may further highlight the capability of repair tools to independently find and gather necessary information. To this end, we collected the top 5 cases with the highest average scores across the five metrics in the No-one-resolve Set, and the 5 cases with the lowest average scores in the All-resolve Set, as shown in Table 3.
@@
110% 213% 147% 166% 2780% 155%
@@
scikit-learn-25747 10 2.5 10 10 0 6.5
seaborn-3407 10 2.5 10 10 0 6.5
django-15202 6.67 5 10 10 0 6.33
scikit-learn-10508 10 5 10 6.67 0 6.33
django-13660 3.33 7.5 10 10 0 6.17
@@
django-12497 3.33 7.5 0 0 0 2.17
django-11133 3.33 2.5 6.67 0 0 2.5
pytest-7432 3.33 2.5 3.33 3.33 0 2.5
pytest-5692 3.33 2.5 6.67 0 0 2.5
django-16046 6.67 2.5 3.33 0 0 2.5
For each case, we examined the execution trajectories generated by seven tools to analyze the reasons they could or could not resolve the respective cases. Due to space limitations, we present only the more commonly observed findings across the seven methods:
scikit-learn-25747: Most methods successfully extracted relevant files, functions, and even fine-grained code snippets containing the error line as a set of suspicious elements for fault localization. This is largely because the issue provided a high-quality description with file- and function-level localization, scoring a 10. However, the main reason for the failure to resolve this case lies in the excessive number of suspicious elements identified by the fault localization module (as each call node in the stack trace tends to be included). The actual error location is deeply embedded within the stack trace, making it challenging to prioritize among the top-ranked elements. This issue tests the model’s ability to infer the relationship between the observed symptoms and the root cause of the error, where the model’s reasoning capabilities remain limited. Additionally, the issue title references a code element associated with the observed error symptom (i.e., the entry point in the stack trace). As the title is a summary of the issue, it naturally draws more attention from the model, further reinforcing the impression that this element caused the error. Meanwhile, in the subsequent repair phase, tools are constrained by the time and cost associated with large-model calls to generate repair suggestions. Resources may be fully exhausted on the first suspicious location, thus limiting opportunities to address subsequent ones.
seaborn-3407: The reason it remains unresolved is nearly identical to that of scikit-learn-25747.
django-15202: The reason it remains unresolved is nearly identical to that of scikit-learn-25747.
scikit-learn-10508: This issue requires consistent modifications at two locations to be resolved. We observed that many tools successfully identified and repaired one location while neglecting the other. In fact, the two required modification points are in close proximity, within the same function and only about 10 lines apart. However, due to current limitations in assessing the completeness of fixes, tools tend to favor generating an ostensibly reasonable solution. Once it passes their designed validation processes—such as using a self-reflective agent to critique the patch’s plausibility or generating plausible reproduction scripts to verify the model-generated patch—the tools often finalize the patch and conclude the resolution process. At present, models demonstrate limited capacity for assessing the completeness of patches in response to issues. Semantic relationships between multiple related modifications that may be overlooked during patch generation are also challenging to detect in subsequent self-reflection phases.
django-13660: The primary reason this issue remains unresolved is due to a misleading solution provided within the issue description. Specifically, the issue description suggests the following fix: “exec should be passed a dictionary containing a minimal set of globals. This can be done by just passing a new, empty dictionary as the second argument of exec.” Many tools adhered to this recommendation, generating patches based on it, primarily modifying “exec(options[’command’])” to “exec(options[’command’], )”, which aligns with the solution suggested in the issue. However, the actual patch replaces the statement with “exec(options[’command’], globals())”, which differs in code semantics from the suggested fix (as shown in the following code box). In the Agentless report, the Quality of Resolve Solution metric received a score of “Complete steps in NL” (7.5 points in Table 4) rather than a “Misleading” score (0 points). We speculate that this may have been an oversight in manual analysis.
Comparison of the Generated Patch Candidate and the Human-written Patch. The Patch Candidate Generated by Gru 1diff --git a/django/core/management/commands/shell.py b/django/core/management/commands/shell.py 2--- a/django/core/management/commands/shell.py 3+++ b/django/core/management/commands/shell.py 4@@ -84,13 +84,13 @@ def python(self, options): 5 def handle(self, **options): 6 # Execute the command and exit. 7 if options[’command’]: 8- exec(options[’command’]) 9+ exec(options[’command’], {}) 10 return 11... The Human-written Patch (Ground Truth) 1diff --git a/django/core/management/commands/shell.py b/django/core/management/commands/shell.py 2--- a/django/core/management/commands/shell.py 3+++ b/django/core/management/commands/shell.py 4@@ -84,13 +84,13 @@ def python(self, options): 5 def handle(self, **options): 6 # Execute the command and exit. 7 if options[’command’]: 8- exec(options[’command’]) 9+ exec(options[’command’], globals()) 10 return 11...
@@
Insights(1)-(2): Overall, we derive two key insights from the above case analyses: (1) Enhancing the large model or agent’s understanding of root causes is crucial. The model’s understanding should extend beyond the location where the issue occurs (i.e., the symptoms) and include deeper reasoning about the relationship between multiple suspicious locations in the issue title or stack trace and the root cause. This would help prioritize the root cause location more effectively. (2) Improving the large model or agent’s ability to generate patches for related locations and to verify patch completeness offers a feasible approach to further enhance issue resolution effectiveness from a holistic perspective.
The above analysis pertains to the five highest-scoring cases within the No-one-resolve Set. Next, we will analyze the five lowest-scoring cases within the All-resolve Set.
django-12497: Although this case’s issue description does not provide any granular bug location information, it does offer a high-quality solution. Repair tools, based on this solution, progressively localized the bug down to the file, function, and line levels, ultimately reusing the suggested solution to generate a correct patch. In relation to the conclusions drawn from Table 3, as the impact of the completeness of these five metrics on issue resolution varies, cases with an equivalent total score but concentrated on more relevant metrics (e.g., line-level location and resolve solution) are likely to be easier to resolve.
django-11133: The patch candidate provided is in a different location within the same file as the manually crafted patch, but it also passes the failed tests. Thus, after the issue provided high-quality file-level bug localization information (6.67 points), the absence of function- and line-level location information does not appear to be critical. However, it is also important to note that passing failed tests does not always equate to generating a semantically correct patch, as discussed in Section  6.
pytest-7432: The patch candidate is situated in a different location within the same function as the manually created patch, but it too passes the failed tests. After providing partial file- and function-level localization information (each scoring 3.33 points), the lack of line-level information does not seem to be a significant limitation.
pytest-5692: Although the issue only provides file-level bug location information, repair tools effectively refined the location through iterative understanding of the issue, eventually pinpointing the bug’s line-level location and producing a correct patch. This indicates that the model and agent possess strong capabilities for identifying relevant information, capable of comprehending the problem described in natural language and extracting high-relevance code segments through multiple rounds of iteration.
django-16046: The reasons for its resolution are nearly identical to those of django-11133.
@@
Insights(3)-(5): Overall, we derive three insights from the above case analyses: (3) Due to the varying impact of the completeness of these five metrics on issue resolution, cases with the same total score but concentrated in more relevant metrics (e.g., line-level location and resolve solution) tend to be easier to resolve. (4) The manually crafted patch may not be the only solution to the issue; multiple resolution strategies may exist. The model may attempt to generate a semantically correct or partially correct patch based on the suspicious locations already highlighted in the issue. (5) Large models and agents demonstrate strong capabilities for discovering relevant information. They can deeply analyze the natural language issue description, iteratively extracting high-relevance code segments over multiple interactions.
As shown in Figure 1, we found significant differences between Agent and non-Agent systems in solving cases. Therefore, we analyzed the differences between cases that can only be solved by the Agent system and those that can only be solved by the non-Agent system. We analyzed two cases that could only be solved by Agentless, specifically django-11422 and sympy-24909. In django-11422, the Agent system identified three locations to make repairs, but bypassed the correct repair location in the golden patch, iter_all_python_module_files(), during the fix. On the other hand, Agentless only selected the location of the golden patch during localization, hence it correctly fixed the issue. The following figure shows the response of the two systems when selecting repair locations.
Comparison of MarsCode Agent and Agentless MarsCode Agent Thank you for providing the iter_all_python_module_files() implementation. It appears that this function only iterates through loaded Python modules, which means manage.py is not included since it’s not imported as a module. Thought: The issue is that manage.py is not being added to the list of watched files. We need to explicitly add it to the extra_files set in the reloader. Let’s modify the runserver command to add manage.py to the watched files. Agentless To fix this, we need to ensure that the manage.py file is included in the list of files to be watched. We can do this by modifying the iter_all_python_module_files function to include the manage.py file explicitly.
In sympy-24909, we also observed that both Agent and non-Agent systems identified the correct faulty location, the function Prefix.__mul__. However, the LLM was not confident about how to fix the bug. Because the Agent system lacked sampling, it directly generated an incorrect patch. On the other hand, Agentless performed 42 samples and chose the patch that appeared most frequently, resulting in the correct modification.
The instance django-12453 is an example that all Agent systems can solve but Agentless cannot. We observed that the Agentless patch is very close to the ground truth, only missing the import of the relevant module. The Agent, through multiple rounds of dialogue and tool invocation, can detect the missing dependencies, whereas Agentless merely generates the code block and cannot compensate for the missing dependencies.
Insights(6)-(8): (6) The Agent system may fail to complete the repair due to an inability to select the correct repair location among multiple candidates. (7) Additionally, the lack of sampling in the Agent system leads to low diversity in the results, causing failure in repairs. (8) Non-Agent systems, on the other hand, may produce syntactically incorrect patches because they lack multi-turn dialogues and tool invocation.

### v2

@@ 4. Analysis & Results
We will sequentially present the analysis results and insights for RQ1 through RQ3.
@@ 4.1. RQ1: Effectiveness of Systems
We analyzed the versions of cases that each of the six tools can individually solve, as well as the differences between the cases that each tool can resolve, as shown in Figure 1. The histogram at the top of the figure shows the number of cases each tool can resolve in the SWE-bench Verified dataset. Specifically, ranked from lowest to highest, these are Emergent E1, Devlo, Learn-by-interact, Midwit Agent, Blackbox AI Agent, W&B Programmer, resolving 286, 291, 301, 311, 314, and 323 cases, respectively. W&B Programmer performs the best, achieving a 13.0% performance improvement over Emergent E1 and addressing 64.6% of the total 500 cases in the SWE-bench Verified dataset. Compared to the popular APR benchmark Defects4J (Just et al., 2014) over the past decade, SWE-bench Verified introduces stricter usage protocols, prohibiting participants from leveraging dynamic evaluation results generated by closely related failing test cases as feedback information for filtering patches. This test case set can only be utilized as a quality standard once the patch generation process has been completed. In this context, considerable of FL methods based on dynamic test execution information (e.g., spectrum-based and mutation-based methods) cannot be used, adding further difficulty to problem detection and resolution. This strict protocol undoubtedly aligns more closely with real-world development scenarios, where repair tools must rely almost solely on issues raised by users and the current state of the code repository to devise solutions. Against this backdrop, W&B Programmer’s ability to address 64.6% of cases underscores its advanced capabilities and utility in real-world development environments.
@@
In Figure 1, the histogram to the right of the tool names presents the count of case versions that different tool combinations can address. Each column indicates the number of case versions solvable by the tools marked with black dots but not by those marked with gray dots. For instance, in the first column, only W&B Programmer is marked with a black dot, while the other five tools are marked with gray dots, indicating that W&B Programmer can solve a unique set of 8 cases that none of the other five tools can handle. Similarly, the seventh column demonstrates that W&B Programmer and Blackbox AI Agent together can solve 3 cases that the other four tools cannot address. The final column shows that there are 181 cases that can be solved by each of 6 tools.
Figure 1. Analysis of Performance of State-of-the-art Techniques on Issue Resolve Task
Analysis of Case Solvability. Among the 500 cases in SWE-bench Verified, 404 cases are solvable by at least one of the six tools, while 96 cases remain unsolved by any tool. Furthermore, 181 cases can be solved by all six tools (represented by the last column in Figure 1). We hypothesize that the issue descriptions for these 404 solvable cases are generally of higher quality, whereas the 96 cases that none of the tools can resolve likely exhibit lower-quality issue descriptions. To validate this hypothesis, we conducted a significance analysis of issue quality differences using five metrics. For quantitative comparison, we assigned a value to each option for each metric, as described in Section 3 (e.g., 3 for StackTrace). The only point to note is that we assigned a score of -1 to the Misleading option in the Quality of Resolve Solutions metric. This ensures that a score of 0 consistently represents "no information" across all metrics. Hence, the scoring range was set from -1 to 3 instead of 0 to 4, making scores across different indicators comparable.
@@
To examine the validity of our hypothesis, we calculated the mean values for the five indicators across two case sets: the set of 96 cases unsolved by any tool (referred to as the "No-one-resolve Set") and the set of 181 cases solvable by all tools (referred to as the "All-resolve Set"). These results are shown in Table 1. For each indicator, a higher score indicates that the corresponding issue description provides more complete and detailed information. It is evident that the scores for all five indicators are consistently higher in the All-resolve Set compared to the No-one-resolve Set. This suggests that issue quality significantly influences the effectiveness of resolution methods, underscoring the importance of crafting clear and comprehensive issue descriptions to improve resolution rates from the outset.
@@
101% 195% 126% 130% 122% 125%
Examining the performance details for each indicator across both case sets, the most notable difference is observed in the Quality of Resolve Solutions metric, where the average score in the All-resolve Set is 1.95 times that of the No-one-resolve Set. This indicates that, at least within the SWE-bench Verified dataset, positive modifications based on observed symptoms can substantially aid in improving the possibility of solving the problem. This is followed by symbol-level, file-level and line-level bug locations, with score differences of 1.30, 1.26 and 1.22 times, respectively. The differences between these three metrics are minimal. Relatively, providing code symbol-level localization results might be more helpful. Lastly, Quality of Reproducible Examples shows only a 1.01-time score difference between the two sets, suggesting that the completeness of this indicator has a comparatively minor impact on enhancing the solvability of bug-fixing systems.
In the above analysis, we validated our hypothesis by calculating differences across five metrics that represent issue quality in the All-resolve Set and No-one-resolve Set, showing a clear disparity in issue quality between cases that can be resolved and those that cannot. This reflects an overall trend, though some outliers may exist. On one hand, we observed several high-scoring cases in the No-one-resolve Set; we aim to investigate why, despite relatively complete information in these issues, no tools were able to resolve them. On the other hand, we also identified some low-scoring cases within the All-resolve Set, which may further highlight the capability of repair tools to independently find and gather necessary information. To this end, we collected the top 5 cases with the highest average scores across the five metrics in the No-one-resolve Set, and the 5 cases with the lowest average scores in the All-resolve Set, as shown in Table 2. It is evident that cases with a relatively high average score but cannot be successfully resolved generally have lower resolve solution quality. Hence, this feature appears to be the most critical factor influencing whether an issue can be resolved, which is consistent with our previous conclusions. For cases with relatively low average scores but can be resolved by all repair systems, their feature distributions vary significantly. Some provide different granular levels of bug location information, while others include reproducible examples and resolve solutions. A case-by-case analysis is needed to draw conclusions.
@@
astropy-13977 3 2 3 3 3 2.8
astropy-14182 3 0 3 3 3 2.4
django-13513 2 3 2 3 2 2.4
django-16667 3 0 3 3 3 2.4
django-16938 3 0 3 3 3 2.4
@@
django-10914 1 0 0 0 1 0.4
sympy-22456 0 0 1 1 0 0.4
django-13516 1 0 0 0 1 0.4
sphinx-8721 1 0 1 0 0 0.4
django-15499 0 2 0 0 0 0.4
We first conducted a manual inspection of the trajectories of the five high-scoring cases that could not be repaired and summarized the following insights:
@@
Insights: (1) Enhancing the understanding of root causes is crucial. The model’s understanding should extend beyond the location where the issue occurs (i.e., the symptoms) and include deeper reasoning about the relationship between multiple suspicious locations in the issue title or stack trace and the root cause. This would help prioritize the root cause location more effectively. (2) Improving the ability to generate patches for related locations and to verify patch completeness offers a feasible approach to further enhance issue resolution effectiveness from a holistic perspective.
Subsequently, we conducted a manual inspection of the trajectories of 5 low-scoring cases that could be repaired by all 6 repair systems, and summarized the following insights:
@@
Insights: (3) The manually crafted patch may not be the only solution to the issue; multiple resolution strategies may exist. The model may attempt to generate a semantically correct or partially correct patch based on the suspicious locations already highlighted in the issue. (4) LLMs and agents demonstrate strong capabilities for discovering relevant information. They can deeply analyze the natural language issue description, iteratively extracting high-relevance code segments over multiple interactions.
As shown in Figure 1, we also discovered that Learn-by-interact (Su et al., 2025) was able to independently repair 12 cases, the highest among the six repair systems. However, there were 30 cases that could be repaired by any of the other five systems, but not by Learn-by-interact, making it the most noteworthy in this aspect.
Learn-by-interact uses an agent that interfaces with offline code repositories to collect extensive, high-quality interaction trajectories. It strategically integrates these historical interactions into prompts via in-context learning, enabling the agent to reference and adapt successful debugging methods to new software issues. This systematic approach results in progressively better bug resolution as experiential knowledge accumulates. In contrast, the other five systems focus on agent process design without leveraging prior experience, which is a relatively novel idea in Learn-by-interact. The significant design differences between the two sets of systems appear to lead to varied impacts on different cases, as the experiment results suggest.
One insight is that generating patch candidates using diverse workflow designs within the process and the voting on the patches via their validity after generation, potentially enhancing the final effectiveness as the number of overall fixable cases increases.
Insights: (5) Using designs that deviate from mainstream processes might help to uncover more repairable cases. For instance, leveraging experience gained from historical repair processes could aid in solving difficult instances. (6) Integrating different repair strategies into the entire workflow to generate a wider variety of patch candidates, and then voting on the validity of these patches, might enhance the final repair effectiveness.
(a) Statistics of samples where at least one buggy file is hit
(b) Statistics of samples where all buggy files are hit
Figure 2. Analysis of Performance of State-of-the-art Techniques on File-level FL Task
(a) Statistics of samples where at least one buggy code symbol is hit
(b) Statistics of samples where all buggy code symbols are hit
Figure 3. Analysis of Performance of State-of-the-art Techniques on Code Symbol-level FL Task
To validate the generalizability of these observations, we further examined ten additional cases. We have to limit the scope of manual analysis due to two primary constraints. First, beyond certain score thresholds, issues no longer met the criteria for being considered extreme, reducing their diagnostic value. Second, each case required in-depth comprehension of the underlying codebase and task context, and the corresponding agent trajectories could span up to 200,000 tokens. This made detailed inspection resource-intensive. To ensure reliability, two authors independently reviewed each case and reconciled their assessments through discussion.